# Gibbs polystability of Fano manifolds, stability thresholds and symmetry breaking

### v2

@@ Abstract.
We extend the probabilistic approach for constructing Kähler–Einstein metrics on log Fano manifolds $(X,\Delta)$ – involving random point processes – to the case of non-discrete automorphism groups, by explicitly breaking the symmetry. This yields a new algebraic notion of Gibbs polystability, conjecturally equivalent to K-polystability. The definition involves a limit of log canonical thresholds on the GIT semistable locus of the $N$-fold products $X^{N},$ that we conjecture coincides with an analytic reduced stability threshold encoding the coercivity of the K-energy functional modulo automorphisms. These conjectures follow from an overarching conjectural Large Deviation Principle for the limit $N\rightarrow\infty.$ We prove several of our conjectures for log Fano curves. By imposing a moment constraint, we derive a strengthened form of the sharp logarithmic Hardy–Littlewood–Sobolev inequality on the two-sphere, that implies the sharp form of Aubin’s refinement of the Moser–Trudinger inequality. In companion papers we will present applications to Onsager’s point vortex model on the two-sphere and the AdS/CFT correspondence.
@@ 1. Introduction
The quest for canonical metrics on complex manifolds is a classical theme in geometry, tracing back to the uniformization theorem. According to the seminal Yau–Tian–Donaldson conjecture a compact complex manifold $X$ admits a Kähler–Einstein metric with positive Ricci curvature, i.e. a Kähler metric with constant positive Ricci curvature, iff $X$ is a K-polystable Fano manifold - this is now a theorem [42, 90]. The Fano condition requires that the anti-canonical line bundle $K_{X}$ of $X$ is ample, while K-polystability is a finer algebro-geometric condition that amounts to a positivity property of all polarized $\mathbb{C}^{*}$–equivariant deformations of $(X,-K_{X}),$ called test configurations. This algebro-geometric characterization has also found striking connections to the AdS/CFT correspondence in theoretical physics, where Kähler–Einstein metrics on Fano manifolds induce Sasaki–Einstein metrics on links that are dual to superconformal gauge theories [82, 44, 43]. More precisely, Fano manifolds yield regular links, while Fano orbifolds give rise to quasi-regular links (and, upon introducing weights, Sasaki–Einstein metrics on general links [6, 76]).
Even though the existence problem for Kähler–Einstein metrics is now settled, it should be emphasized that the problem of explicitly describing the Kähler–Einstein metric in terms of algebro-geometric data of $X$ is, in general, intractable, and only very few explicit examples are known. This problem is one of the motivations for the probabilistic approach to the construction of Kähler–Einstein metrics on $X$ introduced in [12, 13], that was further applied to the AdS/CFT correspondence in [28]. In the case of positive Ricci curvature, it requires, however, that the Fano manifold $X$ does not admit any holomorphic vector fields, i.e. that the group $\text{Aut}_{0}(X)$ of automorphism of $X$ homotopic to the identity, be trivial. In the present work the framework is extended to Fano manifolds with nontrivial $\text{Aut}_{0}(X)$. In companion papers we present applications to some toric quiver gauge theories appearing in the the AdS/CFT correspondence [5] and to Onsager’s statistical mechanical point vortex model on the two-sphere [4].
@@ 1.1. Recap of the case when $\text{Aut}_{0}(X)$ is trivial
Before stating our main results, let us recall the probabilistic approach on a Fano manifold $X$ with trivial group $\text{Aut}_{0}(X),$ introduced in [13]. The starting point is the basic fact that a Kähler–Einstein (KE) metric $\omega_{\text{KE}}$ with positive Ricci curvature on $X$ can be readily be recovered from its (normalized) volume form $dV_{\text{KE}}.$ The strategy of the probabilistic approach is thus to construct the normalized volume form $dV_{KE}$ by sampling a large number of $N$ points on $X.$ More precisely, the goal is to construct a canonical symmetric probability measure on the $N$-fold products $X^{N}$ – solely defined in terms of explicit algebro-geometric data of $X$ – and show that $dV_{\text{KE}}$ emerges as $N\rightarrow\infty,$ when sampling $N$ points. The construction of the probability measure on $X^{N}$ is based on the observation that $N$-fold products $X^{N}$ come with an algebraic measure $\nu^{(N)},$ uniquely determined up to scaling. More precisely, $N$ is the subsequence defined by
@@
However, a necessary condition for $X$ to be Gibbs stable is that $\text{Aut}_{0}(X)$ be trivial. In fact, when $\text{Aut}_{0}(X)$ is non-trivial there is no probability measure on $X^{N}$ that is invariant under the diagonal action of $\text{Aut}_{0}(X)$ (under a mild regularity assumption) [20]. The probabilistic approach in [13] is thus too canonical, in a certain sense, and we will need to break the $\text{Aut}_{0}(X)$–symmetry. This is in line with the fact the group $\text{Aut}_{0}(X)$ acts non-trivially on any given Kähler–Einstein metric. Hence, there is no canonical way of singling out a unique Kähler–Einstein metric.
@@ 1.2. Gibbs polystability and microscopic stability thresholds
We will consider the general framework of log Fano manifolds $(X,\Delta),$ consisting of a complex projective manifold $X$ and an effective $\mathbb{Q}$–divisor $\Delta$ on $X,$ whose log anti-canonical line bundle $K_{(X,\Delta)}$ is an ample $\mathbb{Q}$–line bundle [24]. This setting encompasses, in particular, Fano orbifolds [61]. Kähler–Einstein metrics on $(X,\Delta)$ are, in general, understood as positive currents $\omega_{\text{KE}}$ in $c_{1}\left(-K_{(X,\Delta)}\right)$ with locally bounded potentials, that restrict to genuine Kähler–Einstein metrics on the complement of the support of $\Delta.$ In particular, $\omega_{\text{KE }}$ satisfies the following equation in sense of currents:
@@
We start by introducing a notion of Gibbs polystability, taking the group $\mathcal{G}$ into account. It is formulated in terms of the Zariski open subvariety $(X^{N})_{\text{ss}}$ of $X^{N},$ defined by the semistable locus of $X^{N},$ with respect to the ample $\mathbb{Q}$–line bundle $K_{(X,\Delta)},$ in the sense of Geometric Invariant Theory (GIT) [64]. We will denote by $\Delta_{N}$ the divisor on $X^{N}$ induced by $\Delta$ that is invariant under permutation of the factors of $X^{N}.$ Thus $(X^{N},\Delta_{N})$ is a log Fano manifold. Denote by $\mathcal{D}^{(N)}$ the effective $\mathbb{Q}$-divisor on $X^{N},$ linearly equivalent to $K_{(X^{N},\Delta_{N})},$ cut out by the the $k^{\mathrm{th}}$ root of the section $\det S^{(N)},$ defined by formula 1.3, in terms of a bases in $H^{0}(X,-kK_{(X,\Delta)}).$ To recall the algebraic definition of Gibbs stability, introduced in [13], consider the microscopic stability threshold $\gamma^{(N)}(X,\Delta)$ and its limit $\gamma(X,\Delta)$ defined by
@@
The first equivalence may, perhaps, be too optimistic, in general. But we present arguments motivating the second one. They are based on conjectural Large Deviation Principle that we shall come back to below, linking the invariant $\gamma(X,\Delta)\mathcal{{}^{G}}$ to the reduced analytic stability threshold defined by
@@
where $\mathcal{M}$ denotes the (log) Mabuchi functional on the space of Kähler metrics in $c_{1}\left(-K_{(X,\Delta)}\right),$ whose minimizers are Kähler–Einstein metrics for $(X,\Delta),$ and $(\mathcal{I}-\mathcal{J})^{\mathcal{G}}$ denotes the following $\mathcal{G}$–invariant version of the non-negative functional $\mathcal{I}-\mathcal{J},$ defined as the difference of the classical energy functionals $\mathcal{I}$ and $\mathcal{J}$ in Kähler geometry
@@
(the definition of $\mathcal{I}$ and $\mathcal{J}$ are recalled in Remark 3.1). In the case of a complex curve, the three functionals $\mathcal{I}-\mathcal{J},$ $\mathcal{I}$ and $\mathcal{J}$ all coincide with the square of the Dirichlet norm of the Kähler potential. On a general Kähler manifold they are all comparable and are therefore commonly used as measures of coercivity (or properness) in Kähler geometry [89, 23, 48]; their $\mathcal{G}$–invariant versions appear when studying coercivity modulo the action of $\mathcal{G}$ [95, 48]. In the present context, however, it will be important to work with the functional $\mathcal{I}-\mathcal{J}$ as this choice leads to a thermodynamical interpretation of the corresponding reduced analytic stability threshold $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}$ (see Prop 3.6).
@@
The formula $\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}=2$ implies, in particular, Aubin’s refinement of the Moser–Trudinger inequality on $S^{2}$ [7, 39, 55, 72] and yields a strengthening of the sharp logarithmic Hardy–Littlewood–Sobolev inequality on $S^{2}$ in [37, 38, 9] under a moment constraint, i.e. a strengthening of the following inequality for all probability measures $\mu$ on $S^{2}:$
@@
where $D(\mu|\mu_{0})$ denotes the entropy of $\mu,$ relative to the $SO(3)$-invariant probability measure $\mu_{0}$ on $S^{2},$ and $E(\mu)$ denotes the logarithmic energy of $\mu$ (explicitely expressed in formula 4.3). In fact, combining our formula for $\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}$ with the uniqueness result for the Liouville equation on $S^{2}$ established in [72] yields:
@@
Let $X=\mathbb{P}^{1},$ identified with the unit-sphere $S^{2}$ in $\mathbb{R}^{3},$ and denote by $\mathcal{M}$ the Mabuchi functional on the space of all Kähler forms $\omega$ in $c_{1}(X),$ normalized so that $\mathcal{M}(\omega)=0$ when $\omega$ is proportional to the $SO(3)$-invariant probability measure $\mu_{0}$ on $S^{2}.$ For any $\omega$ in $c_{1}(X),$
$\mathcal{M}(\omega)\geq(\mathcal{I}-\mathcal{J})^{\mathcal{G}}(\omega)$
with equality iff $\omega$ is proportional to $\mu_{0}$ modulo the action of $\mathcal{G}$ (i.e. modulo conformal diffeomorphisms of $S^{2}).$ Moreover, the functional $(\mathcal{I}-\mathcal{J})^{\mathcal{G}}$ may be replaced by $\mathcal{I}-\mathcal{J}$ if the Kähler potential of $\omega$ satisfies Tian’s orthogonality condition 3.31. More generally, for any probability measure $\mu$ on $S^{2},$ satisfying the moment constraint $\int_{S^{2}\subset\mathbb{R}^{3}}\boldsymbol{x}\mu=0,$
@@
The inequality for $\mathcal{M}$ in formula 1.7, for some (non-effective) positive constant $\epsilon$ and $C_{\epsilon}\in\mathbb{R}$ was first established in [48]. Since $\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}=2$ (by Theorem 1.3) the inequality for $\mathcal{M}$ fails on $\mathbb{P}^{1}$ when $\epsilon>1$ (for any $C_{\epsilon})$ and thus the multiplicative constants in Theorem 1.4 are sharp. More generally, the formula for $\delta^{\mathrm{A}}(X,w)^{\mathcal{G}}$ in Theorem 1.3 yields a quasi-sharp generalization of Aubins’ improvement of the Moser-Trudinger inequality to $w\neq 0.$ We show that the discontinuity of $\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}},$ as $w$ decreases to $0,$ is a reflection of a somewhat surprising phenomenon of spontaneous symmetry breaking, appearing when $w\in]0,1/2[$ (see Section 4.1). This is related to the fact that the invariant $\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}$ then differs from the invariant obtained by restricting $\mathcal{M}$ to the subspace of $\mathcal{K}$–invariant Kähler metrics, for a given a maximally compact subgroup $\mathcal{K}$ of $\mathcal{G}$ (and replacing $\mathcal{G}$ with its center), appearing implicitly in [48, 63, 77] – see Section 4.2.
Beyond complex curves, Gibbs stability has turned out to be notoriously difficult to establish, and at present no examples are known apart from the case of exceptional Fano orbifolds, where it holds automatically. But in the companion paper [5] we exhibit non-exceptional log Fano manifolds $(\mathbb{P}^{2},\Delta)$ that are Gibbs polystable already at the minimal level (i.e. for the smallest $k\in]0,\infty[$ such that $kK_{(X,\Delta)}$ defines a line bundle):
Theorem 1.5.
[5] Let $\Delta_{w}$ be the torus-invariant divisor on $\mathbb{P}^{2}$ assigning the same weight $w\in]0,1[$ to each of the three torus-invariant hyperplanes. Then the log Fano manifold $(\mathbb{P}^{2},\Delta_{w})$ is Gibbs polystable at the minimal level iff $w>3/4.$
The log Fano manifold $(\mathbb{P}^{2},\Delta_{w})$ may, when w=11/mw=1-1/m for a positive integer $m,$ be identified with the Fano orbifold defined as the quotient of $\mathbb{P}^{2}$ under the standard action of $(\mathbb{Z}/m\mathbb{Z})^{2}.$ In this case the previous theorem applies to the AdS/CFT correspondence, where the Fano orbifold in question encodes a particular toric quiver gauge theory.

### v3

@@ Abstract.
We extend the probabilistic approach for constructing Kähler–Einstein metrics on log Fano manifolds $(X,\Delta)$ – involving random point processes – to the case of non-discrete automorphism groups, by breaking the symmetry using a moment map constraint. In particular, an algebraic notion of Gibbs polystability is introduced, ensuring that the corresponding point processes on $X$ are well-defined. We conjecture that the Gibbs polystability of $(X,\Delta)$ is equivalent to the existence of a Kähler–Einstein metric and that the unique such metric with vanishing moment emerges when sampling a large number of $N$ points on $X.$ The definition of Gibbs polystability involves a limit of log canonical thresholds on the GIT semistable locus of the $N$-fold products $X^{N},$ that we conjecture coincides – as $N$ tends to infinity – with an analytic reduced stability threshold, encoding the coercivity of the K-energy functional modulo automorphisms. These conjectures follow from an overarching conjectural Large Deviation Principle for the large NN-limit. We prove several of our conjectures on log Fano curves and derive a strengthened form of the sharp logarithmic Hardy–Littlewood–Sobolev (HLS) inequality on the two-sphere, under a moment constraint. It yields quantitative stability results for the sharp logarithmic HLS inequality with optimal stability constants. Furthermore, we show that any log Fano manifold that is strongly uniformly Gibbs polystable admits a Kähler–Einstein metric. In companion papers we will present applications to Onsager’s point vortex model on the two-sphere and the AdS/CFT correspondence.
@@ 1. Introduction
The quest for canonical metrics on complex manifolds is a classical theme in geometry, tracing back to the uniformization theorem. According to the seminal Yau–Tian–Donaldson conjecture a compact complex manifold $X$ admits a Kähler–Einstein metric with positive Ricci curvature, i.e. a Kähler metric with constant positive Ricci curvature, iff $X$ is a K-polystable Fano manifold – this is now a theorem [53, 108]. The Fano condition requires that the anti-canonical line bundle $K_{X}$ of $X$ is ample, while K-polystability is a finer algebro-geometric condition that amounts to a positivity property of all polarized $\mathbb{C}^{*}$–equivariant deformations of $(X,-K_{X}),$ called test configurations. This algebro-geometric characterization has also found striking connections to the AdS/CFT correspondence in theoretical physics, where Kähler–Einstein metrics on Fano manifolds induce Sasaki–Einstein metrics on links that are dual to superconformal gauge theories [100, 57, 56]. More precisely, Fano manifolds yield regular links, while Fano orbifolds give rise to quasi-regular links (and, upon introducing weights, Sasaki–Einstein metrics on general links [7, 94]).
Even though the existence problem for Kähler–Einstein metrics is now settled, it should be emphasized that the problem of explicitly describing the Kähler–Einstein metric in terms of algebro-geometric data of $X$ is, in general, intractable, and only very few explicit examples are known. This problem is one of the motivations for the probabilistic approach to the construction of Kähler–Einstein metrics on $X$ introduced in [13, 14], that was further applied to the AdS/CFT correspondence in [30]. In the case of positive Ricci curvature, it requires, however, that the Fano manifold $X$ does not admit any holomorphic vector fields, i.e. that the group $\text{Aut}_{0}(X)$ of automorphisms of $X$ homotopic to the identity be trivial. In the present work the framework is extended to Fano manifolds with nontrivial $\text{Aut}_{0}(X)$. In companion papers we present applications to some toric quiver gauge theories appearing in the the AdS/CFT correspondence [6] and to Onsager’s statistical mechanical point vortex model on the two-sphere [5].
@@ 1.1. Recap of the case when $\text{Aut}_{0}(X)$ is trivial
Before stating our main results, let us recall the probabilistic approach on a Fano manifold $X$ with trivial group $\text{Aut}_{0}(X),$ introduced in [14]. The starting point is the basic fact that a Kähler–Einstein (KE) metric $\omega_{\text{KE}}$ with positive Ricci curvature on $X$ can be readily be recovered from its (normalized) volume form $dV_{\text{KE}}.$ The strategy of the probabilistic approach is thus to construct the normalized volume form $dV_{KE}$ by sampling a large number of $N$ points on $X.$ More precisely, the goal is to construct a canonical symmetric probability measure on the $N$-fold products $X^{N}$ – solely defined in terms of explicit algebro-geometric data of $X$ – and show that $dV_{\text{KE}}$ emerges as $N\rightarrow\infty,$ when sampling $N$ points. The construction of the probability measure on $X^{N}$ is based on the observation that $N$-fold products $X^{N}$ come with an algebraic measure $\nu^{(N)},$ canonically determined up to scaling. More precisely, $N$ is the subsequence defined by
@@
However, a necessary condition for $X$ to be Gibbs stable is that $\text{Aut}_{0}(X)$ be trivial. In fact, when $\text{Aut}_{0}(X)$ is non-trivial there is no probability measure on $X^{N}$ that is invariant under the diagonal action of $\text{Aut}_{0}(X)$ (under a mild regularity assumption) [21]. The probabilistic approach in [14] is thus too canonical, in a certain sense. This is in line with the fact the group $\text{Aut}_{0}(X)$ acts non-trivially on any given Kähler–Einstein metric – as a consequence there is no canonical way of singling out a unique Kähler–Einstein metric. In order to probabilistically construct a Kähler–Einstein metric we will thus need to break the $\text{Aut}_{0}(X)$–symmetry. This is explained in Section 1.5 below, but we first introduce an algebro-geometric notion of Gibbs polystability, taking the action of the group $\text{Aut}_{0}(X)$ into account and explore the relations to stability thresholds.
@@ 1.2. Gibbs polystability and microscopic stability thresholds
We will consider the general framework of log Fano manifolds $(X,\Delta),$ consisting of a complex projective manifold $X$ and an effective $\mathbb{Q}$–divisor $\Delta$ on $X,$ whose log anti-canonical line bundle $K_{(X,\Delta)}$ is an ample $\mathbb{Q}$–line bundle [26]. This setting encompasses, in particular, Fano orbifolds [78]. Kähler–Einstein metrics on $(X,\Delta)$ are, in general, understood as positive currents $\omega_{\text{KE}}$ in $c_{1}\left(-K_{(X,\Delta)}\right)$ with locally bounded potentials, that restrict to genuine Kähler–Einstein metrics on the complement of the support of $\Delta.$ In particular, $\omega_{\text{KE }}$ satisfies the following equation in the sense of currents:
@@
We start by introducing an algebro-geometric notion of Gibbs polystability, taking the group $\mathcal{G}$ into account. It is formulated in terms of the Zariski open subvariety $(X^{N})_{\text{ss}}$ of $X^{N},$ defined by the semistable locus of $X^{N},$ with respect to the ample $\mathbb{Q}$–line bundle $K_{(X,\Delta)},$ in the sense of Geometric Invariant Theory (GIT) [83]. We will denote by $\Delta_{N}$ the divisor on $X^{N}$ induced by $\Delta$ that is invariant under permutation of the factors of $X^{N}.$ Thus $(X^{N},\Delta_{N})$ is a log Fano manifold. Denote by $\mathcal{D}^{(N)}$ the effective $\mathbb{Q}$-divisor on $X^{N},$ linearly equivalent to $K_{(X^{N},\Delta_{N})},$ cut out by the the $k^{\mathrm{th}}$ root of the section $\det S^{(N)},$ defined by formula 1.3, in terms of a bases in $H^{0}(X,-kK_{(X,\Delta)}).$ To recall the algebraic definition of Gibbs stability, introduced in [14], consider the microscopic stability threshold $\gamma^{(N)}(X,\Delta)$ and its limit $\gamma(X,\Delta)$ defined by
@@
The first equivalence may, perhaps, be too optimistic, in general. But we present arguments motivating the second one. They are based on a conjectural Large Deviation Principle that we shall come back to below, linking the invariant $\gamma(X,\Delta)\mathcal{{}^{G}}$ to the reduced analytic stability threshold defined by
@@
where $\mathcal{M}$ denotes the (log) Mabuchi functional on the space of Kähler metrics in $c_{1}\left(-K_{(X,\Delta)}\right),$ whose minimizers are Kähler–Einstein metrics for $(X,\Delta),$ and $(\mathcal{I}-\mathcal{J})^{\mathcal{G}}$ denotes the following $\mathcal{G}$–invariant version of the classical non-negative functional $\mathcal{I}-\mathcal{J},$ defined as the difference of the classical energy functionals $\mathcal{I}$ and $\mathcal{J}$ in Kähler geometry
@@
(the definition of $\mathcal{I}$ and $\mathcal{J}$ are recalled in Remark 3.1. In the case of a complex curve, all the three functionals $\mathcal{I}-\mathcal{J},$ $\mathcal{I}$ and $\mathcal{J}$ are proportional to the Dirichlet norm of the Kähler potential. On a general Kähler manifold they are all comparable and are therefore commonly used as measures of coercivity (or properness) in Kähler geometry [107, 25, 61]; their $\mathcal{G}$–invariant versions appear when studying coercivity modulo the action of $\mathcal{G}$ [113, 61]. In the present context, however, it will be important to work with the functional $\mathcal{I}-\mathcal{J}$ as this choice leads to a thermodynamical interpretation of the corresponding reduced analytic stability threshold $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}}$ (see Prop 3.6). It should be emphasized that, while the definition of $\mathcal{I}-\mathcal{J}$ depends on the choice of a reference Kähler metric $\omega_{0}\in c_{1}(X)$, this does not affect the definition 1.7 of $\delta^{\mathrm{A}}(X,\Delta)^{\mathcal{G}},$ since it only alters the additive constant $C_{\epsilon}.$
@@
The formula for $\delta^{\mathrm{A}}(X,w)^{\mathcal{G}}$ in Theorem 1.3 yields a generalization of Aubins’ improvement of the Moser–Trudinger inequality to $w\neq 0$ with a quasi-sharp multiplicative constant. We show that the discontinuity of $\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}},$ as $w$ decreases to $0,$ is a reflection of a somewhat surprising phenomenon of spontaneous symmetry breaking, appearing when $w\in]0,1/2[$ (see Section 4.1). This is related to the fact that the invariant $\delta^{\mathrm{A}}(X,\Delta_{w})^{\mathcal{G}}$ then differs from the invariant obtained by restricting $\mathcal{M}$ to the subspace of $\mathcal{K}$–invariant Kähler metrics, for a given a maximally compact subgroup $\mathcal{K}$ of $\mathcal{G}$ (and replacing $\mathcal{G}$ with its center), appearing implicitly in [61, 81, 95] – see Section 4.2.
Coming back to Conjecture 1.2for a general log Fano manifold, we introduce a slight modification of the log canonical threshold $\gamma^{(N)}(X,\Delta)^{\mathcal{G}},$ denoted by $\gamma_{\epsilon_{N}}^{(N)}(X,\Delta)^{\mathcal{G}},$ obtained by replacing the semi-stable locus $(X^{N})_{\text{ss}}$ by a Zariski open subset $(X^{N})_{\text{ss},\epsilon_{N}}$ of $X^{N}$ containing slightly unstable configurations (see Definition 7.1). This leads to a notion of strong uniform Gibbs polystability, defined by replacing $(X^{N})_{\text{ss}}$ with $(X^{N})_{\text{ss},\epsilon_{N}}$ (which is expected to be equivalent to uniform Gibbs polystability, as illustrated by the case of log Fano curves; see Example 7.2). We establish the following quantitative lower bound on the reduced analytic stability threshold $\delta^{A}(X,\Delta)^{\mathcal{G}}:$
Theorem 1.4.
Let $(X,\Delta)$ be a log Fano manifold whose Futaki character vanishes. Then there exists a constant $C$ such that
$(1-Ck^{-1})\gamma_{\epsilon_{N}}^{(N)}(X,\Delta)^{\mathcal{G}}\leq\delta^{A}(X,\Delta)^{\mathcal{G}}.$
Hence, if $(X,\Delta)$ is strongly uniformly Gibbs polystable, then $(X,\Delta)$ admits a Kähler–Einstein metric and, as consequence, $(X,\Delta)$ is K-polystable.
The previous theorem generalizes [73, Cor 2.6], where it was shown – using purely algebro-geometric techniques – that $\gamma^{(N)}(X,\Delta)$ is bounded from above by Fujita–Odaka’s invariant $\delta_{k}(X,\Delta)$ converging to $\delta(X,\Delta)$ and, as a consequence,
$\limsup_{N\rightarrow\infty}\gamma^{(N)}(X,\Delta)\leq\delta(X,\Delta)$
(which can be made quantiative, using that $\delta_{k}(X,\Delta)(1-O(1/k))=\delta(X,\Delta)$, by [55]). However, it seems challenging to extend the algebro-geometric argument in [73] so as to incorporate the action of the group $\mathcal{G}.$ Instead, our proof of Theorem 1.4 adapts the analytic approach developed in [20], by proving an effective large deviation bound (Theorem 7.4), motivated by the conjectural Large Deviation Principle alluded to above.
1.3. Higher dimensional examples
Beyond complex curves, Gibbs stability has turned out to be notoriously difficult to establish, except for the case of exceptional log Fano manifolds, where it holds automatically at any level $k.$ In Section 2.5.1 we provide the first non-exceptional higher-dimensional examples of Gibbs stable log Fano varieties. More precisely, we show that the Fano orbifold defined as the quotient of the Fermat hypersurface $X_{n+1}$ of degree $(n+1)$ in $\mathbb{P}^{n+1},$ under the standard action of the group $(\mathbb{Z}/\mathbb{Z}_{n})^{n},$ is Gibbs stable at any level $k.$ This Fano orbifold may be identified with a log pair $(\mathbb{P}^{n},\Delta),$ where $\Delta$ is a divisor that assigns the same weight 11/(n+2)1-1/(n+2) to $n+2$ normally crossing hyperplanes in $\mathbb{P}^{n}.$ Although the group $\text{Aut}_{0}(\mathbb{P}^{n},\Delta)$ is trivial, the proof of the Gibbs stability of $(\mathbb{P}^{n},\Delta)$ exploits that the group $\text{Aut}_{0}(\mathbb{P}^{n})$ is non-trivial. Furthermore, in the companion paper [6] we exhibit log Fano manifolds $(\mathbb{P}^{2},\Delta)$ with non-trivial group $\text{Aut}_{0}(\mathbb{P}^{n},\Delta)$ that are Gibbs polystable at the minimal level (i.e. for the smallest $k\in]0,\infty[$ such that $kK_{(X,\Delta)}$ defines a line bundle):
Theorem 1.5.
[6] Let $\Delta_{w}$ be the torus-invariant divisor on $\mathbb{P}^{2}$ assigning the same weight $w\in[0,1[$ to each of the three torus-invariant hyperplanes. Then the log Fano manifold $(\mathbb{P}^{2},\Delta_{w})$ is Gibbs polystable at the minimal level iff $w>3/4$ or $w=0$.
The conclusion that $\mathbb{P}^{2}$ is Gibbs polystable at the minimal level, but $(\mathbb{P}^{2},\Delta_{w})$ is not Gibbs polystable for $w>0$ sufficiently small shows that the discontinuity phenomenon observed for $(\mathbb{P}^{1},\Delta_{w})$ in Theorem 1.3 persists in higher dimensions, at least for this example and at the minimal level. The case of $w=0$ in Theorem 1.5 is a special case of Proposition 2.16, stating that $\mathbb{P}^{n}$ is Gibbs polystable at the minimal level for any $n$. When w=11/mw=1-1/m for a positive integer $m,$ the log Fano manifold $(\mathbb{P}^{2},\Delta_{w})$ may be identified with the Fano orbifold defined as the quotient of $\mathbb{P}^{2}$ under the standard action of $(\mathbb{Z}/m\mathbb{Z})^{2}.$ In this case the previous theorem applies to the AdS/CFT correspondence, where the Fano orbifold in question encodes a particular toric quiver gauge theory.
1.4. A sharp constrained log HLS inequality on $S^{2}$ and application to quantitative stability
@@
Using the formula $\delta^{\mathrm{A}}(\mathbb{P}^{1},0)^{\mathcal{G}}=2$ in Theorem 1.3 we establish a strengthening of the sharp logarithmic Hardy–Littlewood–Sobolev inequality on $S^{2}$ in [43, 44, 10] under a moment constraint, i.e. a strengthening of the following inequality for all probability measures $\mu$ on $S^{2}:$
@@
where $D(\mu|\mu_{0})$ denotes the entropy of $\mu,$ relative to the $\mathrm{SO}(3)$–invariant probability measure $\mu_{0}$ on $S^{2}$ (aka the Kullback–Leibler divergence) and $E(\mu)$ denotes the logarithmic energy of $\mu$ (explicitely expressed in formula 4.3). More precisely, combining our formula for $\delta^{\mathrm{A}}(\mathbb{P}^{1},0)^{\mathcal{G}}$ with the uniqueness result for the Liouville equation on $S^{2}$ established in [80] leads – via a somewhat miraculous optimization identity (formula 5.3), established in the course of the proof – to the following theorem:
@@
Let $X=\mathbb{P}^{1},$ identified with the unit-sphere $S^{2}$ in $\mathbb{R}^{3},$ and denote by $\mathcal{M}$ the Mabuchi functional on the space of all Kähler forms $\omega$ in $c_{1}(X),$ normalized so that $\mathcal{M}(\omega)=0$ when $\omega$ is proportional to the $\mathrm{SO}(3)$–invariant probability measure $\mu_{0}$ on $S^{2}.$ For any $\omega$ in $c_{1}(X),$
(1.10) $\mathcal{M}(\omega)\geq(\mathcal{I}-\mathcal{J})^{\mathcal{G}}(\omega)$
with equality iff $\omega$ is proportional to $\mu_{0}$ modulo the action of $\mathcal{G}$ (i.e. modulo conformal diffeomorphisms of $S^{2}).$ Moreover, the functional $(\mathcal{I}-\mathcal{J})^{\mathcal{G}}$ may be replaced by $\mathcal{I}-\mathcal{J}$ if the Kähler potential of $\omega$ satisfies Tian’s orthogonality condition 3.30. More generally, for any probability measure $\mu$ on $S^{2},$ satisfying the moment constraint $\int_{S^{2}\subset\mathbb{R}^{3}}\boldsymbol{x}\mu=0,$
@@
The inequality for $\mathcal{M}$ in formula 1.7 on a general log Fano manifold for some (non-effective) positive constant $\epsilon$ and $C_{\epsilon}\in\mathbb{R}$ was first established in [61]. Since $\delta^{\mathrm{A}}(X,0)^{\mathcal{G}}=2$ (by Theorem 1.3) the inequality for $\mathcal{M}$ fails on $\mathbb{P}^{1}$ when $\epsilon>1$ (for any $C_{\epsilon})$ and thus the multiplicative constants in Theorem 1.6 are sharp. Exploiting that there is no additive constant in inequality 1.10 we deduce the following quantitative stability results for the sharp logarithmic HLS inequality 1.9 on $S^{2}$ with an optimal stability constant $1/2$ (see also Cor 5.7 for a more general formulation):
Corollary 1.7.
The following quantitative stability inequality for the sharp logarithmic HLS inequality on $S^{2}$ holds
(1.12) $D(\mu|\mu_{0})-E(\mu)\geq\frac{1}{2}\inf_{\nu:\,D(\nu)=E(\nu)}D\left(\mu|\nu\right),$
for any probability measure $\mu$ on $S^{2}$ such that $D(\mu|\mu_{0})<\infty.$ Moreover, the stability constant $1/2$ is optimal, i.e. the stability inequality above fails when $1/2$ is replaced by any strictly larger constant.
The inf in the right hand side of the stability inequality 1.12 can be interpreted as an entropic “distance” (or more precisely divergence) between $\mu$ and the three-dimensional submanifold of $\mathcal{P}(S^{2})$ consisting of the optimizers of the sharp HLS inequality. As explained in Section 5.2 the previous corollary appears to be the first instance where the optimal stability constant has been computed explicitly, in the context of fractional Sobolev and HLS inequalities or their logarithmic versions. It improves, in particular, on the recent quantitative stability result in [47].